Solve the following exact differential equation with initial value. (5x + 4y)dx + (4x - 8y3)dy = 0, y(0) = 2
Answer #1

Given differential equation is (5x + 4y) dx + ( u x-843) dy=0 (0)=2 which is of the form of Moxt Noy =o.. where M=5x 4 yy. 3वत्र+ d y=const. not cont. JS - - - - - - 5x + 4xy - 899 = - - 6 + 4xy – 24 5x + 4xy - =C given y (o) - 2 0+0-32 = C =-32 the